OpenStudy (anonymous):

solve for y (y/4) - (y-1)/2 ≥ 3 a. y ≤ -10 b. y ≥ -10 c. y ≤ 10 d. y ≥ 10

OpenStudy (anonymous):

a. from Mathematica.\[\text{Reduce}\left[\frac{y}{4}-\frac{y-1}{2}\geq 3\right] \]\[y\leq -10 \]

OpenStudy (anonymous):

thank you:)